40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4020     Entry time: Tue Dec 7 16:09:53 2010
Author: josephb 
Type: Update 
Category: CDS 
Subject: c1iscex status 

I swapped out the IO chassis which could only handle 3 PCIe cards with the another chassis which has space for 17, but which previously had timing issues.  A new cable going between the timing slave and the rear board seems to have fixed the timing issues. 

I'm hoping to get a replacement PCI extension board which can handle more than 3 cards this week from Rolf and then eventually put it in the Y-end rack.  I'm also still waiting for a repaired Host interface board to come in for that as well.

At this point, RFM is working to c1iscex, but I'm still debugging the binary outputs to the analog filters.  As of this time they are not working properly (turning the digital filters on and off seems to have no effect on the transfer function measured from an excitation in SUSPOS, all the way around to IN1 of the sensor inputs (but before measuring the digital fitlers).  Ideally I should see a difference when I switch the digital filters on and off (since the analog ones should also switch on and off), but I do not.

ELOG V3.1.3-